OpenStudy (anonymous):

The standard form of a Linear Equation is Ax + By = C. On another page, transform this into the slope-intercept form using A,B,C,x, and y.

OpenStudy (gorv):

slope intercept from is y=mx+c so ax+by=c subtract ax from both side by=c-ax by=-ax+c divide both side by b y=-ax/b+c/b